21g of sodium carbonate is dissolved in water and the solution made up to 500cm cube. Find the molarity of the solution. Naco3
1
Expert's answer
2017-02-24T14:42:06-0500
C = n_subst. / V_sol. (moll/L)

500 cm 3 = 500 ml = 0.5 L

M(Na 2 CO 3 ) = 2×23 + 12 + 3×16 = 106 g/moll

n(Na 2 CO 3 ) = 21/106 = 0.2 moll

C(Na 2 CO 3 ) = 0.2/0.5 = 0.4 moll/L

Answer: 0.4 moll/L (Na 2 CO 3 )